OpenStudy (anonymous):

70% of the people in a town are adult. Find the number of children if the total population of the town is 35000

OpenStudy (anonymous):

is 500

OpenStudy (anonymous):

Maybe .7 * 35000 ??? Just a wild guess... :-)

OpenStudy (anonymous):

you divide 7 in 35000 is 500

OpenStudy (anonymous):

i agree^

OpenStudy (anonymous):

Know you don't divide by 7 70% is 70/100, so you multiply 35000 by .7 (that's POINT 7) And it is NOT 500.

OpenStudy (anonymous):

no. of adults=35000*70/100=24500 no. of children=35000-24500=10500

OpenStudy (anonymous):

Another no-brainer way of looking at this: Is 70% MORE THAN HALF? Is 500 MORE THAN HALF of 35000? Obviously, 500 is NOT the answer...

OpenStudy (anonymous):

And one more thing, just to keep kicking the horse... 35000/7 isn't 500 anyway...it's 5000

OpenStudy (anonymous):

I'm done... :-)

OpenStudy (anonymous):

Qweqwe is actually wrong. Luis' answer is the correct one. A more simple way to approach this is that if 70% of the total population are adults, 30% are children. All you would need to do is 0.3 * 3500 This also is equal to 10500

OpenStudy (anonymous):

I meant 35000. 0.3 * 35000 Missed a zero :P

OpenStudy (anonymous):

g